# Disparate Impact on Group Accuracy of Linearization for Private Inference

### v1

@@ 1 Introduction
Figure 1: ResNet18 accuracy on UTKFaces across various ReLU linearization budgets using SNL Cho et al. (2022). 100% ReLUs corresponds to the original model. Subgroup sample sizes are shown above the corresponding bars. The accuracy for the majority group remains almost unchanged while other groups, in particular the minority one, are diversely impacted.
@@ 2 Related Work
Piece-wise approximation with ReLU functions. In addition to observing disparate impacts from methods that reduce the number of ReLU functions in neural networks, our work seeks to explain these effects by linking them to the expressiveness and approximation capabilities of a Deep Neural Network (DNN). These studies trace back to seminal works in Cybenko (1989); Hornik (1991). The widespread use of ReLU activations to introduce non-linearity in the learning of decision boundaries has motivated early studies about their approximation power Montúfar et al. (2014); Pan and Srikumar (2016). Several papers have observed the effect of ReLU activations in the learning process Hanin and Rolnick (2019); Lin and Jegelka (2018); Lu et al. (2017); Shen et al. (2022); Yarotsky (2017). Among others, Arora et al. (2018); Liu and Liang (2021) distinguish themselves for contributions to the field, analyzing the connection between the number of ReLU activations and that of the approximated piecewise functions, along with a ReLU-dependent bound on the approximation error.
Finally, model compression has also been shown to induce fairness issues. Empirical observations reported that quantization, network compression, and knowledge distillation could amplify the unfairness in different learning tasks Lukasik et al. (2022); Hooker et al. (2020a, b); Joseph et al. (2020); Blakeney et al. (2021); Ahn et al. (2022).
Our paper takes the cue from these works to show how ReLU linearization techniques may adversely affect the fairness of the resulting predictors.
@@ 3 Settings
4 Why ReLU reduction may increase unfairness?
This section introduces theoretical insights to elucidate the causes behind the observed disparity effects resulting from the linearization of ReLU reduction. We start by recalling some fundamental results that will help us develop our analysis in the rest of the paper. The first results Liu and Liang (2021) define a connection between the number of piecewise functions and the approximation error of convex functions. Throughout the section, ${\cal F}$ denotes the set of strict convex univariate functions.
@@ 4 Why ReLU reduction may increase unfairness?
The second result Arora et al. (2018) upper-bounds the number of piecewise linear functions that can be represented by a ReLU network.
@@
Figure 2: Decision boundary estimated by a ReLU model (left) and a linearized model (right). The ground truth decision boundary (black dotted line) separate majority (blue) from minority (red) samples. The linearized model shows an inferior approximation of the decision boundary, and therefore a lower prediction accuracy. Moreover, its decision boundary is deeper into to the minority class resulting in accuracy disparities.
@@ 4.1 Decision Boundary Approximation and Fairness
We note that the bound in Theorem 4.4 depends also on the distance $\|\bm{\tilde{\theta}}-\bm{\theta}\|$ between the ReLU reduced model and the original model. Generally, it is not trivial to establish a formal relationship between such distance and the rate of ReLU linearization; indeed such a formalization depends inherently on the model and the ReLU distribution induced by the linearization scheme adopted. However, our empirical analysis provides a strong sign for such a trend to occur, as it will be further elaborated in Section 5.2, Figure 7.
@@ 5 Empirical Results on the Fairness Analysis
We next present the key findings from our empirical analysis on how ReLU linearization affects the fairness of linearized models. These findings build upon and extend the theoretical insights from earlier sections to scenarios involving high-dimensional and non-convex decision boundaries. In summary, we show that (1) model linearization produces disparate degradation of performance across groups and that such disparity is enhanced with the larger amounts of linearization; (2) the relative magnitude of gradient norms for a group serves as a reliable indicator of the group’s accuracy decline resulting from ReLU linearization; (3) groups located farther from (or closer to) the decision boundary are less (more) likely to experience accuracy reduction at varying levels of ReLU linearization.
@@
  • CIFAR-10 Krizhevsky and Hinton (2009). Containing $60,000$ $32\!\times\!32$ RGB images of 10 classes (airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks).
To assess the effect of network linearization on groups with various representations, we evaluate several models and architectures on unbalanced datasets. While UTKFace and SVHN are naturally unbalanced, we unbalanced CIFAR-10 by selecting 90% of the samples for each of 5 classes, and retaining 10% of the samples for each of the remaining classes. All reported metrics are average over 10 random seeds. The UTKFace dataset includes images with a large variation in age, ranging from 0 to 116 years old. As customary in demography, the ages are categorized into four groups: 2-12 (young children), 13-17 (adolescents), 18-64 (working-age individuals), and over 64 (senior citizens) for the experiments in this paper and these groups contain, respectively, $10.14\%$, $3.63\%$, $77.56\%$, and $8.67\%$ of the total samples in the dataset. Data points corresponding to ages 0 and 1 are considered outliers and thus omitted for this task. Race classification on UTKFace utilizes the dataset’s provided labels: white, black, Asian, native Indian, and others which correspond to $42.51\%$, $19.09\%$, $14.49\%$, $16.77\%$, and $7.14\%$ of the total samples in the dataset, respectively. Finally, digit recognition studies are conducted on the naturally imbalanced SVHN dataset. This dataset contains 10 digits from 0 to 9 which have $6.75\%$, $18.92\%$, $14.45\%$, $11.60\%$, $10.18\%$, $9.39\%$, $7.82\%$, $7.64\%$, $6.89\%$, and $6.36\%$ of the total samples in the dataset, respectively.
@@
Linearization methods. The paper considers two ReLU linearization techniques, representing the current state of the art: Selective Network Linearization (SNL) Cho et al. (2022) and DeepReDuce (DR) Jha et al. (2021) [^1] We also attempted to incorporate results from SENet Kundu et al. (2023), but faced challenges. Although a validation code for the published models is available, a public implementation of the linearization algorithm is missing. Despite our efforts, we could not obtain the code or reproduce the results. Given that SENet aims to preserve global accuracy, we suspect it may encounter fairness issues similar to other frameworks considered.. Both methods aim to linearize ReLU s in neural networks, but their approaches differ: SNL selectively linearizes specific ReLU s, whereas DR provides for a much coarser linearization strategy. In particular, DR does not allow to control the “distribution” of the ReLU functions across the architecture, but only to linearize blocks of contiguous nodes. This fundamental difference is expected to result in distinct patterns of disparity observed with each method. As a matter of fact, DR is not inherently data-driven, and its performance is not only affected by the number of residual ReLU functions, but also by the block that the practitioner decides to linearize. In turn, this results in different linearized model configurations as reported in Figure 4(b).
In the following, the term “ReLU budget" is used to refer to the number of ReLU s retained in a ReLU-linearized model across both methods. However, note that specifying a desired ReLU budget is feasible only with SNL. In contrast, with DR, the number of retained ReLUs depends on the specific layers linearized. Additionally, while ReLU linearization may lead to a loss in utility, notice that the models and ReLU budgets used for the analysis yield a minor decrease in global test accuracies (as shown in Figure 5) indicating that these configurations represent practical, deployable ReLU-linearized models.
@@ 5.1 Impact on Accuracy
Recall that Theorem 4.4 highlighted the connection between a drop in loss (as a differentiable proxy for accuracy) experienced by a given group, under a certain ReLU linearization level, and two key characteristics: the gradient norms associated with such groups at convergence and the distance to the decision boundary. While these results hold for a restricted class of functions, the next two sections provide further evidence of such relationships on highly non-linear models for a variety of linearization techniques and architectures.
@@ 5.3 Distance to the Decision Boundary and Fairness
This observation, in conjunction with what is observed in Figure 7 (left), underscores the key relation between distance to the decision boundary plays and fairness issues following ReLU linearization. The understanding of this interplay is crucial for the design of the proposed mitigation technique, reviewed in the next section.
@@ 6 Mitigation and impact on fairness
So far, we have presented an in-depth analysis of the disparate effects that linearization techniques produce on the fairness of unbalanced datasets, showing how reduced ReLU budgets negatively affect the accuracy over underrepresented groups. We will now present a solution to extend ReLU linearization to contexts where fairness matters along with Private Inference. Recall from Proposition 4.5 the relationship between a group’s Hessian eigenvalues, their proximity to the decision boundary, and the corresponding errors made by the model for that group. Since, as shown in the previous section, the distance to the decision boundary of a sample directly relates to their accuracy drop across various ReLU linearization rates, our strategy leverages this intuition and introduces “fairness regularizers” to equalize the distances to the decision boundary of various groups. These regularizers, capture the differences in losses between groups and population, (and by proxy the distance to the boundary). The method leverages Lagrangian duality principles Fioretto et al. (2020) for implementation and it works off-the-shelf with the Knowledge-Distillation (KD) based finetuning step, offering a seamless integration into existing algorithms.
@@
        $\rho_{\bm{\tilde{\theta}}^{r},\bm{\theta},B}\leftarrow|\mathcal{L}^{\text{KD}}(f_{\bm{\tilde{\theta}}^{r}},f_{\bm{\theta}},B)-\mathcal{L}^{\text{KD}}_{a}(f_{\bm{\tilde{\theta}}^{r}},f_{\bm{\theta}},B)|$
        $\bm{\tilde{\theta}}^{r}\leftarrow\bm{\tilde{\theta}}^{r}-\alpha\nabla_{\bm{\tilde{\theta}}^{r}}\left(\mathcal{L}(f_{\bm{\tilde{\theta}}^{r}},f_{\bm{\theta}},B)+\bm{\lambda}_{e}^{\top}\rho_{\bm{\tilde{\theta}}^{r},\bm{\theta},B}\right)$
@@
     ${\bm{\lambda}_{e+1}\leftarrow\bm{\lambda}_{e}+\mu|\mathcal{L}^{\text{KD}}(f_{\bm{\tilde{\theta}}^{r}},f_{\bm{\theta}},S_{T})-\mathcal{L}^{\text{KD}}_{a}(f_{\bm{\tilde{\theta}}^{r}},f_{\bm{\theta}},S_{T})|}$
@@
Figure 8 illustrates the effect of the mitigation strategy on the UTKFace dataset for both SNL (left) and DR (right) methods, with results segmented by age (top) and race (bottom) labels. Negative values indicate accuracy improvements for a group compared to the baseline accuracy of the original model at the same level of ReLU reduction. This visualization brings to light two key aspects: First, the mitigation consistently enhances accuracy across all groups and linearization levels tested, with minimal negative effects on majority classes. Second, notice how the minority groups improve their relative performance, a trend that is especially evident for DR at higher levels of ReLU linearization.
Furthermore, as detailed in Section C.3, there’s an observable reduction in gradient norms and an augmentation in the distance to the decision boundary for minority groups under the mitigation approach. These results are important: they illustrate the potential for effective mitigations to significantly improve the performance of minority groups without detriment to majority groups, thus substantially reducing the unfairness instigated by ReLU linearization.

### v3

@@ Abstract
marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the conf style. Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
Saswat Das* 1  Marco Romanelli* 2  Ferdinando Fioretto1
footnotetext: *Equal contribution 1University of Virginia, Charlottesville, VA, USA 2New York University, New York, NY, USA. Correspondence to: Saswat Das <saswatdas@email.virginia.edu>, Marco Romanelli <mr6852@nyu.edu>.
@@ 1 Introduction
Figure 1: ResNet18 accuracy on UTKFace across various ReLU linearization budgets using SNL Cho et al. (2022). 100% ReLUs corresponds to the original model. Subgroup sample sizes are shown above the corresponding bars. The accuracy for the majority group remains almost unchanged while other groups, in particular the minority one, are diversely impacted.
@@ 2 Related Work
Piece-wise Approximation with ReLU Functions. In addition to observing disparate impacts from methods that reduce the number of ReLU functions in neural networks, our work seeks to explain these effects by linking them to the expressiveness and approximation capabilities of a Deep Neural Network (DNN). These studies trace back to seminal works in Cybenko (1989); Hornik (1991). The widespread use of ReLU activations to introduce non-linearity in the learning of decision boundaries has motivated early studies about their approximation power Montúfar et al. (2014); Pan & Srikumar (2016). Several papers have observed the effect of ReLU activations in the learning process Hanin & Rolnick (2019); Lin & Jegelka (2018); Lu et al. (2017); Shen et al. (2022); Yarotsky (2017). Among others, Arora et al. (2018); Liu & Liang (2021) distinguish themselves for contributions to the field, analyzing the connection between the number of ReLU activations and that of the approximated piecewise functions, along with a ReLU-dependent bound on the approximation error.
Privacy and Fairness Tradeoff in Machine Learning. Among other works on the tradeoff between privacy and fairness in machine learning, Bagdasaryan et al. (2019) provides mostly empirical evidence of the disparate effect of differentially private training algorithms on fairness, while Farrand et al. (2020) further investigates this phenomenon by considering different levels of differential privacy and unbalance in the datasets. These works provide empirical evidence to show the noise injection and the gradient clipping involved in differentially private Stochastic Gradient Descent impact the overall accuracy, and especially that of underrepresented classes. The privacy framework we consider is extremely different, as the diverse impact is not a direct consequence of the privacy preservation framework design, but of the architectural adjustments made to speed up cryptographic secure inference . Such adjustments impact the approximation capabilities of the model resulting in disparate accuracy across different groups.
Finally, model compression has also been shown to induce fairness issues. Empirical observations reported that quantization, network compression, and knowledge distillation could amplify the unfairness in different learning tasks Lukasik et al. (2022); Hooker et al. (2020a; b); Joseph et al. (2020); Blakeney et al. (2021); Ahn et al. (2022). Inspired by these works, we show how ReLU linearization techniques may adversely affect the fairness of the resulting predictors.
@@ 3 Settings
4 Why Network Linearization May Increase Unfairness?
This section introduces theoretical insights to elucidate the causes behind the observed disparity effects resulting from the linearization of ReLU reduction. We start by recalling some fundamental results that will help us develop our analysis in the rest of the paper. The first results Liu & Liang (2021) defines a connection between the number of piecewise functions and the approximation error of convex functions. Throughout the section, ${\cal F}$ denotes the set of strict convex univariate functions.
@@ 4 Why Network Linearization May Increase Unfairness?
The second result Arora et al. (2018) upper-bounds the number of pieacewise linear functions that can be represented by a ReLU network.
@@
Figure 2: Decision boundary estimated by a ReLU model (left) and a linearized model (right). The ground truth decision boundary (black dotted line) separates majority (blue) from minority (red) samples. The linearized model shows an inferior approximation of the decision boundary, and therefore a lower prediction accuracy. Moreover, its decision boundary is deeper into to the minority class resulting in accuracy disparities.
@@ 4.1 Decision Boundary Approximation and Fairness
$\blacksquare$ Remark about Theorem 4.4. It is important to note that the bound presented in Theorem 4.4 also depends on the distance $|\bm{\tilde{\theta}}-\bm{\theta}|$ between the ReLU-linearized model and the original model. Generally, a formal relationship between this distance and the rate of ReLU linearization is challenging to establish, as it inherently depends on the specific model and the distribution of ReLUs induced by the adopted linearization scheme.
Empirically, our experiments consistently observe this result, as shown in Figure 8 (left): the distance $|\bm{\tilde{\theta}}-\bm{\theta}|$ decreases with the increase of the ReLU budget. It is important to note that the model parameters retain information about the activation functions and the data split $S_{a}$ used during the training phase when the gradients of the loss w.r.t. the parameters are computed. Thus, a weight vector $\bm{\tilde{\theta}}$ may induce different outputs based on where in the network the various ReLU activations are preserved. Additionally, while our results are based on networks that have been extensively trained, possibly to or very close to convergence, certain corner cases could affect our analysis of the bound. For instance, the bound may not hold for poorly trained networks or if we were to impose equal weights on linearized and non linearized models. In that case the distance $\|\bm{\tilde{\theta}}-\bm{\theta}\|$ would reduce the bound to zero, while the left-hand side of the bound would still be positive by construction.
To be more precise, the function $J(\cdot)$ should be considered not solely as a function of the model’s parameters, but also of the data, and the set of activation functions. This would lead to a more complex bound that increases with the distance between the channels, i.e., the posterior probabilities induced by $h_{\bm{\theta}}(\cdot)$ and $h_{\bm{\tilde{\theta}}}(\cdot)$ on $S_{a}$. While this refinement could render our analysis more precise, it would also introduce much notation, while also not altering our key observation that linearizing models may have disparate impacts on underrepresented groups. $\blacksquare$
@@ 5 Empirical Results on the Fairness Analysis
We next present the key findings from our empirical analysis on how ReLU linearization affects the fairness of linearized models. These findings build upon and extend the theoretical insights from earlier sections to scenarios involving high-dimensional and non-convex decision boundaries. In summary, we show that (1) model linearization produces disparate degradation of performance across groups and that such disparity is enhanced with the larger amounts of linearization; (2) the relative magnitude of gradient norms for a group serves as a reliable indicator of the group’s accuracy decline resulting from ReLU linearization; (3) groups located farther from (or closer to) the decision boundary are less (more) likely to experience accuracy reduction at varying levels of ReLU linearization [^1] Code: https://github.com/SaswatD27/ICML_Linearization_Disparate_Impact.
@@
  • CIFAR-10 Krizhevsky & Hinton (2009). Containing $60,000$ $32\!\times\!32$ RGB images of 10 classes (airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks).
To assess the effect of network linearization on groups with various representations, we evaluate several models and architectures on unbalanced datasets. While UTKFace and SVHN are naturally unbalanced, we unbalanced CIFAR-10 by selecting 90% of the samples for each of 5 classes, and retaining 10% of the samples for each of the remaining classes. All reported metrics are average over 10 random seeds. The UTKFace dataset includes images with a large variation in age, ranging from 0 to 116 years old. As customary in demography, the ages are categorized into four groups: 2-12 (young children), 13-17 (adolescents), 18-64 (working-age individuals), and over 64 (senior citizens) for the experiments in this paper and these groups contain, respectively, $10.14\%$, $3.63\%$, $77.56\%$, and $8.67\%$ of the total samples in the dataset. Data points corresponding to ages 0 and 1 are considered outliers and thus omitted for this task.Race classification on UTKFace utilizes the dataset’s provided labels: white, black, Asian, native Indian, and others which correspond to $42.51\%$, $19.09\%$, $14.49\%$, $16.77\%$, and $7.14\%$ of the total samples in the dataset, respectively. Finally, digit recognition studies are conducted on the naturally imbalanced SVHN dataset. This dataset contains 10 digits from 0 to 9 which have $6.75\%$, $18.92\%$, $14.45\%$, $11.60\%$, $10.18\%$, $9.39\%$, $7.82\%$, $7.64\%$, $6.89\%$, and $6.36\%$ of the total samples in the dataset, respectively
@@
Linearization Methods. The paper considers two ReLU linearization techniques, representing the current state of the art: Selective Network Linearization (SNL) Cho et al. (2022) and DeepReDuce (DR) Jha et al. (2021). Both methods aim to linearize ReLU s in neural networks, but their approaches differ: SNL selectively linearizes specific ReLU s, whereas DR provides for a much coarser linearization strategy. In particular, DR does not allow to control the “distribution” of the ReLU functions across the architecture, but only to linearize blocks of contiguous nodes. This fundamental difference is expected to result in distinct patterns of disparity observed with each method. As a matter of fact, DR is not inherently data-driven, and its performance is not only affected by the number of residual ReLU functions, but also by the block that the practitioner decides to linearize. In turn, this results in different linearized model configurations as reported in Figure 5(b) [^2] We also attempted to incorporate results from SENet Kundu et al. (2023), but faced challenges. Although a validation code for the published models is available, a public implementation of the linearization algorithm is missing. Despite our efforts, we could not obtain the code or reproduce the results. Given that SENet aims to preserve global accuracy, we suspect it may encounter fairness issues similar to other frameworks considered..
In the following, the term “ReLU budget” is used to refer to the number of ReLU s retained in a ReLU-linearized model across both methods. However, note that specifying a desired ReLU budget is feasible only with SNL. In contrast, with DR, the number of retained ReLUs depends on the specific layers linearized. Additionally, while ReLU linearization may lead to a loss in utility, notice that the models and ReLU budgets used for the analysis yield a minor decrease in global test accuracies (as shown in Figure 5) indicating that these configurations represent practical, deployable ReLU-linearized models.
@@ 5.1 Impact on Accuracy
Recall that Theorem 4.4 highlighted the connection between a drop in loss (as a differentiable proxy for accuracy) experienced by a given group, under a certain ReLU linearization level, and two key characteristics: the gradient norms associated with such groups at convergence and the distance to the decision boundary. While these results hold for a restricted classes of functions, the next two sections provide further evidence on such relationships on highly non linear models for a variety of linearization techniques and architectures. Further experiments on different architectures and wider variants of the ResNet models are reported in Figure 7.
(a) SNL on UTKFace (race) for VGG-19.
(b) SNL on UTKFace (age) for VGG-19.
(c) SNL on UTKFace (race) for WideResNet.
(d) SNL on UTKFace (age) for WideResNet.
Figure 7: Results for VGG-19 and WideResNet models on UTKFace with SNL. We observe that the performance is consistent with the results obtained for the ResNet18 and ResNet34 models: more linearization produces higher accuracy loss for underrepresented groups, while the accuracy remains almost unchanged in the majority group. We observe that the phenomenon of disparity in the accuracy for underrepresented groups is evident on WideResNet-22-8, a wider variant of ResNet.
@@ 5.3 Distance to the Decision Boundary and Fairness
This observation, in conjunction with what observed in Figure 8 (left), underscores the key relation between distance to the decision boundary plays and fairness issues following ReLU linearization. Understanding of this interplay is crucial to design the proposed mitigation technique, reviewed in the next section.
@@ 6 Mitigation and Impact on Fairness
So far, we have presented an in-depth analysis of the disparate effects that linearization techniques produce on the fairness of unbalanced datasets, showing how reduced ReLU budgets negatively affect the accuracy over underreprested groups. We will now present a solution to extend ReLU linearization to contexts where fairness matters along with Private Inference. Recall from Proposition 4.5 the relationship between a group’s Hessian eigenvalues, their proximity to the decision boundary, and the corresponding errors made by the model for that group. Since, as shown in the previous section, the distance to the decision boundary of a sample directly relates to their accuracy drop across various ReLU linearizaiton rates, our strategy leverages this intuition and introduces “fairness regularizers” to equalize the distances to the decision boundary of various groups. These regularizers, capture the differences in losses between groups and population, (and by proxy the distance to the boundary). The method leverages Lagrangian duality principles Fioretto et al. (2020) for implementation and it works off-the-shelf with the Knowledge-Distillation (KD) based finetuning step, offering a seamless integration into existing algorithms.
@@
        $\rho_{\bm{\tilde{\theta}}^{r},\bm{\theta},B}\leftarrow|\mathcal{L}^{\text{KD}}(h_{\bm{\tilde{\theta}}^{r}},h_{\bm{\theta}},B)-\mathcal{L}^{\text{KD}}_{a}(h_{\bm{\tilde{\theta}}^{r}},h_{\bm{\theta}},B)|$
        $\bm{\tilde{\theta}}^{r}\leftarrow\bm{\tilde{\theta}}^{r}-\alpha\nabla_{\bm{\tilde{\theta}}^{r}}\left(\mathcal{L}(h_{\bm{\tilde{\theta}}^{r}},h_{\bm{\theta}},B)+\bm{\lambda}_{e}^{\top}\rho_{\bm{\tilde{\theta}}^{r},\bm{\theta},B}\right)$
@@
     ${\bm{\lambda}_{e+1}\leftarrow\bm{\lambda}_{e}+\mu|\mathcal{L}^{\text{KD}}(h_{\bm{\tilde{\theta}}^{r}},h_{\bm{\theta}},S_{T})-\mathcal{L}^{\text{KD}}_{a}(h_{\bm{\tilde{\theta}}^{r}},h_{\bm{\theta}},S_{T})|}$
@@
Figure 9 illustrates the effect of the mitigation strategy on the UTKFace dataset for both SNL (left) and DR (right) methods, with results segmented by age (top) and race (bottom) labels. Negative values indicate accuracy improvements for a group compared to the baseline accuracy of the original model at the same level of ReLU reduction. There are two key aspects: First, the mitigation consistently enhances accuracy across all groups and linearization levels tested, with minimal negative effects on majority classes. Second, notice how the minority groups improve their relative performance, a trend that is especially evident for DR at higher levels of ReLU linearization. Furthermore, there is an observable reduction in gradient norms and an augmentation in the distance to the decision boundary for minority groups under the mitigation approach (see Section D.1). These results illustrate the potential for effective mitigations to significantly improve the performance for minority groups without detriment to majority groups, thus substantially reducing the unfairness produced by ReLU linearization.
Finally, an ablation study on the impact of the parameter $\mu$ in the mitigation algorithm is presented in Section D.2.